OpenStudy (anonymous):

Given the function g(x) = 4x − 5, compare and contrast g(2) and g(−4). Choose the statement that is true concerning these two values. The value of g(2) is larger than the value of g(−4). The value of g(2) is smaller than the value of g(−4). The value of g(2) is the same as the value of g(−4). The values of g(2) and g(−4) cannot be compared.

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

g(2)=4(2)-5=3, g(-4)= -16-5=-21. so g(2) is larger. This would be answer A

OpenStudy (anonymous):

g(2)= 3 g(-4)= -21

OpenStudy (anonymous):

Compare them now.

OpenStudy (anonymous):

3 is larger

OpenStudy (anonymous):

Yes g(2) will be larget than g(-4)